16/the cubed root of 16
so sixteen divided by the cubed root of sixteen
thanks i would really appericiate it. i have been working on it for about 2 hrs now.
sorry wrong email the first time its capl3147@student.tamut.edu
capl is the first four letters of my last name caple

Answer by frankiej(1)

16/ cubrt(16) = 16/ cubrt(2*2*2*2) = 16/cubrt(2^3*2)= 16/(2)[(cubrt2)] =
8/cubrt(2)